# SvANet: A Scale-variant Attention-based Network for Small Medical Object Segmentation

### v3

@@ III-C Cross-scale Feature Guidance
III-D Scale-variant Attention
The concept of cross-scale feature guidance is based on convolution operations, which have inherent limitations in processing global feature representations. While global pooling operations can facilitate learning context representations, it is restricted to handling features in a uniform scale manner. Given a subregion $x_{j}$ of an input tensor $x$, the output of vanilla global attention, denoted as $\mathcal{A}(x)$, is calculated as follows:
@@ III-D Scale-variant Attention
Figure 3: Output feature maps from cross-scale guidance without or with scale-variant attention (SvAttn). (a) Input image and (b) ground truth. The feature maps are generated without using SvAttn in (c) $s=1,t=4$ and (d) $s=3,t=4$, and with using SvAttn in (e) $s=1,t=4$ and (f) $s=3,t=4$. Example images in odd and even rows include ultra-small and small medical objects, respectively.
@@
As indicated by Eq. 4, the subregions located at the same proportional scaling position across stages are dynamic. This variability enables the cross-scale guidance module to effectively discern the relationships between the high-level and low-level features. Consequently, SvAttn enhances the network’s capability to recognize downsized small medical objects throughout a sequence of stages. As illustrated in Fig. 3ce and df, for the same source and target stages, the features captured using SvAttn are more detailed and comprehensive for both ultra-small and small medical objects compared to those obtained without using SvAttn. For example, from top to bottom, there is a more precise delineation of networked retinal vessels, more discernible morphology of nevi, more pronounced instance boundaries of organs such as kidneys, and finer details in sperm morphology. In contrast, without using SvAttn, critical features such as retinal vessels of glaucoma in the first and second rows, the nevus in the third row, and the kidneys and cyst in the sixth rows were overlooked. It is noteworthy that ultral-small objects are harder to perceive compared to small objects without using SvAttn. For example, moving downwards from the odd rows of Fig. 3cd, no retinal vessel were discovered, a relatively small nevus region was highlighted, and the left kidney was missed.
@@ III-E Convolution with Vision Transformer
The proposed convolution with vision transformer by assembling tensors, termed AssemFormer, is illustrated in top-right dased green boxes in Fig. 1. Inspired by [33, 16], AssemFormer incorporates a 3×3 convolution and a 1×1 convolution, followed by two transformer blocks and two convolution operations. AssemFormer bridges convolution and transformer operations by stacking and unstacking feature maps. Equipped with this design, AssemFormer tackles the issue of lacking inductive biases for the vanilla transformer.
@@
Furthermore, a skip connection and concatenation are incorporated to mitigate the information loss concerning small medical objects. Leveraging the convolution-transformer hybrid structure, the AssemFormer block can simultaneously learn the local and global representations of an input medical image. According to the ablation study presented in the fifth and sixth rows of Tab. IV, the AssemFormer significantly improves the segmentation performance of SvANet.
@@ IV-A1 Dataset
The FIVES dataset comprises 800 fundus photographs taken with ophthalmoscopes featuring age-related macular degeneration, diabetic retinopathy, glaucoma, and health fundus types. The ISIC 2018 dataset includes skin lesion images collected by dermatoscopes, encompassing both healthy and unhealthy skin areas. PolypGen, sourced from six different hospitals using colonoscopes, focuses exclusively on polyps. The ATLAS dataset consists of 90 MRI scans of livers, detailing two types of medical objects: the liver and tumor. KiTS23 has 599 CT scans of kidneys, categorized into three semantic classes (i.e., kidney, tumor, and cyst). For experimental comparisons, the ATLAS and KiTS23 datasets were sliced into sequences of 2-D images. In addition, TissueNet includes images of cells from the pancreas, breast, tonsil, colon, lymph, lung, esophagus, skin, and spleen, derived from humans, mice, and macaques, collected using cell imaging platforms such as CODEX and CyCIF, with annotations for whole cells and their nuclei.
@@ IV-A2 Implementation Details and Evaluation Metric
In this study, the mini-batch size was set to 4. Data augmentation strategies applied to pre-process the input images included random horizontal flips, random cropping to a resolution of 512×512, Gaussian blur, distortion, and rotation. The AdamW optimizer [35] and a cross-entropy loss function were utilized, with the learning rate decaying from 5×10-5 to 1×10-6 following a cosine schedule [36]. The total training process spanned 100 epochs. The results were calculated by averaging the outcomes from three times of training and testing cycles.
@@ IV-B Results for Datasets with Diverse Object Sizes
As presented in Tab. II, SvANet outperformed other SOTA methods across three object scales in the FIVES, ISIC 2018, and ATLAS datasets, excluding sensitivity of 93.54% & 87.13% in ISIC 2018 & ATLAS datasets and MAE of 5.35×10-4 & 0.66×10-4 in FIVES & ATLAS datasets for the “all” object scale, as summarized in Tab. II. Additionally, SvANet surpasses other methods with increments in the mDice of +2.95% & +5.23%, mIoU of +2.36% & +5.78%, sensitivity of +0.19% & +5.03%, and F2 score of +1.28% & +5.15% for differentiating ultra-small and small retinal vessels in the FIVES dataset. However, the MAE is comparatively high ($>$75×10-4) in ultra-small retinal vasculature segmentation across all tested models, potentially due to the minimal number (4) ultra-small objects providing insufficient learnable features for deep learning algorithms. In ISIC 2018 and ATLAS datasets, SvANet excelled in segmenting ultra-small objects (i.e., skin lesions, livers, and hepatic tumors) with mDice of 96.11% & 89.79%, mIoU of 92.76% & 86.06%, sensitivity of 98.35% & 86.68%, and F2 score of 97.42% & 87.71%. These results suggest significant potential for SvANet in diagnosing dermatological skin lesions and hepatic tumors in MRI scans, particularly for objects with an area ratio smaller than 1% or 10%. Thus, SvANet can ameliorate therapeutic approaches such as excision therapy, laser therapy, electrosurgery, and radiotherapy for treating these conditions.
Furthermore, the segmentation results for the PolypGen and KiTS23 datasets demonstrate that SvANet delivers superior performance than other SOTA methods across three object scales. Specifically, SvANet achieved the highest mDice of 84.15% & 96.12%, 91.17% & 94.01%, and 93.16% & 94.54% for ultra-small, small, and all medical object scales in PolypGen and KiTS23 datasets, respectively. Moreover, SvANet delivered up to 14.83% & 1.22%, 6.23% & 2.80%, and 6.93% & 2.70% increments in F2 score over other tested methods for ultra-small, small, and all object scales in PolypGen and KiTS23 datasets, separately. The F2 score, the harmonic mean of sensitivity and precision, underscores the robustness of SvANet in medical object segmentation. SvANet also recorded the lowest MAE, 1.01×10-4 & 0.02×10-4, 0.66×10-4 & 0.07×10-4, and 0.81×10-4 & 0.08×10-4 across three object scales for PolypGen and KiTS23 datasets, indicating a high level of precision in the pixel-level recognition of polyps, kidneys, renal tumors, and cysts.
@@
ISIC 2018 UNet (MICCAI’15) [9] 87.22 89.82 91.10 81.14 82.75 83.93 62.99 11.28 6.91 97.49 93.75 90.64 91.92 92.05 90.82
@@
In the TissueNet dataset, which includes only ultra-small and small cells, Tab. II reveals that the SvANet leads in segmentation performance, achieving 80.25% & 88.05% in mDice, 71.60% & 79.45% in mIoU, 7.22×10-4 & 3.28×10-4 in MAE, 83.36% & 88.07% in sensitivity, and 82.00% & 88.06% in F2 score, across ultra-small and small medical object scales, respectively. Notably, SvANet performance is essentially distinguished in the segmentation of ultra-small tissue cells, surpassing other SOTA models by at least 9.60% in mIoU, 8.50% in mDice, and 6.61% in F2 score. This superior performance contrasts with improvements of less than 5% observed in the five other datasets, as shown in Tab. II, which may be attributed to the relatively large number of ultra-small objects in TissuNet (i.e., 9096 cells, approximately ten times more objects than other datasets).
Furthermore, the mDice results trends for all tested methods across ultra-small, small, and all medical object segmentation in FIVES, ISIC 2018, PolypGen, ATLAS, KiTS23, and TissueNet datasets are illustrated in Fig. 4. This figure highlights the SvANet, represented by the red line, consistently leads across diverse object scales and datasets. In the FIVES dataset, as shown in Fig. 4a, only SvANet exhibits an increasing mDice as object scale increases, while other methods’ mDice initially increases and then decreases. The subbranches of retinal vessels are relatively thin and the number of vessels increases as the occupied area expands. Therefore, the subbranches become more difficult to be discriminated, decreasing mDice as the object scale range expands from $\leq$10to $\leq$100%.
Figure 5: Examples of segmentation results in the ATLAS, KiTS23, and SpermHealth datasets in different rows for error analysis. Examples contain ultra-small objects (i.e., tumor, cyst, and sperm), small objects (i.e., kidney), and objects with $>$10% area ratio (i.e., liver). (a) Input image, (b) Ground truth, and prediction results of (c) HRNet, (d) TransNetR, and (e) SvAttn (ours).
However, SvANet maintains a growing trend without decline, demonstrating its effectiveness in recognizing retinal vasculatures, which is crucial for diagnosing blindness-causing diseases. Besides, in ISIC 2018 and KiTS23 datasets, SvANet and over half of other methods exhibit a mDice trend resembling an “L” shape, as depicted in Fig. 4be. Fewer ultra-small objects in these datasets introduce significant variability, likely contributing to this “L” trend. In the PolypGen, ATLAS, and TissueNet datasets, there is a consistent increase in mDice trends, as shown in Fig. 4cdf. Notably, no change is observed in TissueNet between the small and all object scales, as both categories contain identical medical images. Closer inspection of the Fig. 4d and Tab. II reveals that SvANet is the only method that achieved a “V” trend in the ATLAS dataset, with the best mDice of 89.79% for segmenting ultra-small compared to small and all sizes of livers and tumors, underscoring SvANet’s capability to effectively discriminate ultra-small medical objects. Additionally, SvANet consistently displays narrow error bars (shown as color bands) across three object scales, indicating its robustness in accurately recognizing medical objects.
IV-C Results for the Dataset for Only Ultra-small Objects
To further evaluate the performance of SvANet in distinguishing ultra-small medical objects, experiments were conducted on the SpermHealth dataset, which exclusively has sperms with an area ratio of less than 1%. As shown in Tab. III, SvANet secured top performance in sperm segmentation within the SpermHealth dataset, achieving 72.58% in mDice, 61.44% in mIoU, 13.06×10-4 in MAE, 72.50% in sensitivity, and 72.51% F2 score. SvANet’s performance in sperm segmentation notably exceeded that of other models, surpassing them by up to 36.66% in mDice, 26.16% in mIoU, 36.76% in sensitivity, and 36.74% in F2 score. Besides, the performance metrics (mDice, mIoU, sensitivity, and F2 score) gained in the SpermHealth dataset are significantly lower than those observed in ISIC 2018, PolypGen, ATLAS, and KiTS23 for all tested models, with a gap of $>$10%, because all sperms have an area lower than 1%, presenting limited learnable features and posing more significant challenges for differentiation.
TABLE IV: Ablation study results in the SpermHealth dataset on the ultra-small category. The best results are underlined in bold.

### v4

@@ III-B Cross-Scale Feature Guidance
III-C Scale-Variant Attention
Cross-scale feature guidance is based on convolution operations, which have inherent limitations in processing global feature representations. While global pooling operations can facilitate learning context representations, it is restricted to handling features in a uniform scale manner. Given a subregion $x_{j}$ of an input tensor $x$, the output of vanilla global attention, denoted as $\mathcal{A}(x)$, is calculated as follows:
@@ III-C Scale-Variant Attention
As indicated by Eq. 3, the subregions located at the same proportional scaling position across stages are dynamic. This variability enables the cross-scale guidance module to effectively discern the relationships between the high-level and low-level features. Consequently, SvAttn enhances the network’s capability to recognize downsized small medical objects throughout a sequence of stages. As illustrated in Fig. 3ce and df, for the same source and target stages, the features captured using SvAttn are more detailed and comprehensive for both ultra-small and small medical objects compared to those obtained without using SvAttn. For example, from top to bottom, there is a more precise delineation of networked retinal vessels, more discernible morphology of nevi, more pronounced instance boundaries of organs such as kidneys, and finer details in sperm morphology. In contrast, without using SvAttn, critical features such as retinal vessels of glaucoma in the first and second rows, the nevus in the third row, and the kidneys and cyst in the sixth rows were overlooked. It is noteworthy that ultra-small objects are harder to perceive compared to small objects without using SvAttn. For example, moving downwards from the odd rows of Fig. 3cd, no retinal vessel was discovered, a relatively small nevus region was highlighted, and the left kidney was missed.
Figure 4: Output feature maps from the MCBottleneck without or with an attention mechanism. Example images in odd and even rows include ultra-small and small medical objects, respectively (GT: ground truth).
III-D Monte Carlo Attention
The Monte Carlo attention (MCAttn) module, as presented by the purple block in Fig. 2, uses a random-sampling-based pooling operation to generate scale-agnostic attention maps, enabling the network to capture relevant information across different scales, enhancing its ability to identify small medical objects. The MCAttn generates attention maps by randomly selecting a 1×1 attention map from three scales: 3×3, 2×2, and 1×1 (pooled tensors). In conventional methods such as squeeze-excitation (SE), global-average pooling is used to acquire a 1×1 output tensor, which helps calibrate the inter-dependencies between channels [21]. However, this approach has limited capacity to exploit cross-scale correlations. To address this limitation, MCAttn calculates the attention maps from features across three scales, thereby enhancing long-range semantic interdependences. Given an input tensor, $x$, the output attention map of MCAttn, denoted as $\mathcal{A}_{\text{m}}(x)$, is computed as follows:
$\mathcal{A}_{\text{m}}(x)=\sum_{i=1}^{n}P_{1}(x,i)f(x,i)$ (6)
The Monte Carlo sampling method described in Eq. 6 allows for the random selection of association probabilities, enabling the extraction of both local information (e.g., angle, edge, and color) and context information (e.g., whole image texture, spatial correlation, and color distribution). In Fig. 4c, the second to the fourth rows and the final row illustrate that MCBottleneck, without using an attention mechanism, struggles to detect the retinas and nevi and often overlooks several sperms. Conversely, when attention mechanisms like SE, CBAM, and CoordAttn are used, localization of densely occupied regions (e.g., optic disc, kidneys, and sperms) is enhanced compared to when no attention mechanism is used. However, sparse regions, such as retinal vasculatures and nevus centers, are often underlooked, especially the ultra-small ones, as shown in Fig. 4c-f. Instead, using MCAttn in MCBottleneck, as depicted in Fig. 4cg, enhances the discernibility of the morphology and precise location of both ultra-small and small medical objects compared to when MCAttn is not used. For instance, in Fig. 4, moving downwards, MCBottleneck coupled with MCAttn emphasizes more apparent retinal vessels for glaucoma, sharper boundaries of nevi, and more perceptible morphology of kidneys, cysts, and sperms. MCAttn also accentuates other medical objects of interest, such as retinas, nevi, kidneys, and sperms, as shown in Fig. 4g.
@@ III-E Convolution with Vision Transformer
The proposed assembly-based convolutional vision transformer, termed AssemFormer, is illustrated in top-right dased green boxes in Fig. 2. Inspired by [34, 18], AssemFormer incorporates a 3×3 convolution and a 1×1 convolution, followed by two transformer blocks and two convolution operations. AssemFormer bridges convolution and transformer operations by stacking and unstacking feature maps. Equipped with this design, AssemFormer tackles the lack of inductive biases for the vanilla transformer.
Figure 5: Output feature maps from the AssemFormer. The feature maps are extracted from four encoder stages individually (GT: ground truth). Since the layer in the fifth stage is directly connected with the ASPP module, no AssemFormer is used at this stage.
@@
Furthermore, a skip connection and concatenation are incorporated to mitigate the information loss concerning small medical objects. Leveraging the convolution-transformer hybrid structure, the AssemFormer block can simultaneously learn the local and global representation of an input medical image. According to the ablation study presented in the fifth and sixth rows of Tab. IV, the AssemFormer significantly improves the segmentation performance of SvANet.
In Fig. 5, progressing from left to right, the AssemFormer increasingly highlights smaller areas that more accurately align with the ground truth, especially notable in scenarios with fewer medical objects. For instance, the first row of Fig. 4 demonstrates how the thin lines of retinal vasculature and light reflections are initially emphasized, becoming progressively thicker. Subsequently, these lines become shorter and more focused on a smaller region corresponding to the optic disc location, as depicted in the second first of Fig. 4ef. Large-scale distortions, such as noise or compression defects, play a role in this observed trend where the concentration of feature maps intensifies with a deeper layer. The pattern of increased feature map concentration is consistent across the segmentation of various medical objects, including skin lesions, polyps, hepatic tumors, livers, kidneys, tissue cells, and sperms.
@@ IV-A1 Dataset
The FIVES dataset comprises 800 fundus photographs taken with ophthalmoscopes featuring age-related macular degeneration, diabetic retinopathy, glaucoma, and health fundus types. The ISIC 2018 dataset includes skin lesion images collected by dermatoscopes, encompassing healthy and unhealthy skin areas. PolypGen, sourced from six different hospitals using colonoscopes, focuses exclusively on polyps. The ATLAS dataset consists of 90 MRI scans of livers, detailing two types of medical objects: the liver and tumor. KiTS23 has 599 CT scans of kidneys, categorized into three semantic classes (i.e., kidney, tumor, and cyst). For experimental comparisons, the ATLAS and KiTS23 datasets were sliced into sequences of 2-D images. In addition, TissueNet includes images of cells from the pancreas, breast, tonsil, colon, lymph, lung, esophagus, skin, and spleen, derived from humans, mice, and macaques, collected using cell imaging platforms such as CODEX and CyCIF, with annotations for whole cells and their nuclei.
@@ IV-A2 Implementation Details and Evaluation Metric
In this study, the mini-batch size was set to 4. Data augmentation strategies applied to pre-process the input images included random horizontal flips, random cropping to a resolution of 512×512, Gaussian blur, distortion, and rotation. The AdamW optimizer [36] and a cross-entropy loss function were utilized, with the learning rate decaying from 5×10-5 to 1×10-6 following a cosine schedule [37]. The total training process spanned 100 epochs. The results were calculated by averaging the outcomes from three times of training and testing cycles. All backbone was pretrained in the ImageNet-1K [38] dataset. Additionally, all tested methods followed the configurations above of training except that nnUNet utilized official settings [13] for training.
@@ IV-B Results for Datasets with Diverse Object Sizes
As presented in Tab. II, SvANet outperformed other SOTA methods across three object scales in the FIVES, ISIC 2018, and ATLAS datasets, excluding sensitivity of 93.54% & 87.13% in ISIC 2018 & ATLAS datasets and MAE of 5.35×10-4 & 6.6×10-3 in FIVES & ATLAS datasets for the “all” object scale, as summarized in Tab. II. Additionally, SvANet surpasses other methods with increments in the mDice of at least +2.95% & +5.23%, mIoU of +1.97% & +5.78%, sensitivity of +0.19% & +5.03%, and F2 score of +1.28% & +5.15% for differentiating ultra-small and small retinal vessels in the FIVES dataset. However, the MAE is comparatively high ($>$7.5×10-3) in ultra-small retinal vasculature segmentation across all tested models, potentially due to the minimal number (4) ultra-small objects providing insufficient learnable features for deep learning algorithms. In ISIC 2018 and ATLAS datasets, SvANet excelled in segmenting ultra-small objects (i.e., skin lesions, livers, and hepatic tumors) with mDice of 96.11% & 89.79%, mIoU of 92.76% & 86.06%, sensitivity of 98.35% & 86.68%, and F2 score of 97.42% & 87.71%. These results suggest significant potential for SvANet in diagnosing dermatological skin lesions and hepatic tumors in MRI scans, particularly for objects with an area ratio smaller than 1% or 10%. Thus, SvANet can ameliorate therapeutic approaches such as excision therapy, laser therapy, electrosurgery, and radiotherapy for treating these conditions.
Furthermore, the segmentation results for the PolypGen and KiTS23 datasets demonstrate that SvANet delivers superior performance than other SOTA methods across three object scales. Specifically, SvANet achieved the highest mDice of 84.15% & 96.12%, 91.17% & 94.01%, and 93.16% & 94.54% for ultra-small, small, and all medical object scales in PolypGen and KiTS23 datasets, respectively. Moreover, SvANet delivered up to 14.83% & 2.76%, 6.23% & 6.88%, and 6.93% & 6.33% increments in F2 score over other tested methods for ultra-small, small, and all object scales in PolypGen and KiTS23 datasets, separately. The F2 score, the harmonic mean of sensitivity and precision, underscores the robustness of SvANet in medical object segmentation. SvANet also recorded the lowest MAE, 1.01×10-4 & 2.0×10-2, 6.6×10-3 & 7.0×10-2, and 8.1×10-3 & 8.0×10-2 across three object scales for PolypGen and KiTS23 datasets, indicating a high level of precision in the pixel-level recognition of polyps, kidneys, renal tumors, and cysts.
@@
ISIC 2018 UNet (MICCAI’15) [10] 88.00 89.77 90.87 80.98 82.59 83.55 47.51 11.22 7.13 98.23 93.67 90.61 93.35 92.00 90.71
@@
In the TissueNet dataset, which includes only ultra-small and small cells, Tab. II reveals that the SvANet leads in segmentation performance, achieving 80.25% & 88.05% in mDice, 71.60% & 79.45% in mIoU, 7.22×10-4 & 3.28×10-4 in MAE, 83.36% & 88.07% in sensitivity, and 82.00% & 88.06% in F2 score, across ultra-small and small medical object scales, respectively. Notably, SvANet performance is essentially distinguished in the segmentation of ultra-small tissue cells, surpassing other SOTA models by at least +9.60% in mIoU, +8.50% in mDice, and +6.61% in F2 score. This superior performance contrasts with improvements of less than 5% observed in the five other datasets, as shown in Tab. II, which may be attributed to the relatively large number of ultra-small objects in TissuNet (i.e., 9096 cells, approximately ten times more objects than other datasets).
Furthermore, the mDice results trends for all tested methods across ultra-small, small, and all medical object segmentation in FIVES, ISIC 2018, PolypGen, ATLAS, KiTS23, and TissueNet datasets are illustrated in Fig. 6. This figure highlights the SvANet, represented by the red line, consistently leads across diverse object scales and datasets. In the FIVES dataset, as shown in Fig. 6a, only SvANet exhibits an increasing mDice as object scale increases, while other methods’ mDice initially increases and then decreases. The subbranches of retinal vessels are relatively thin, and the number of vessels increases as the occupied area expands. Therefore, the subbranches become more difficult to discriminate, decreasing mDice as the object scale range expands from $\leq$10to $\leq$100%. However, SvANet maintains a growing trend without decline, demonstrating its effectiveness in recognizing retinal vasculatures, which is crucial for diagnosing blindness-causing diseases. Besides, in ISIC 2018 and KiTS23 datasets, SvANet and over half of other methods exhibit a mDice trend resembling an “L” shape, as depicted in Fig. 6be. Fewer ultra-small objects in these datasets introduce significant variability, likely contributing to this “L” trend. In the PolypGen, ATLAS, and TissueNet datasets, there is a consistent increase in mDice trends, as shown in Fig. 6cdf. Notably, no change is observed in TissueNet between the small and all object scales, as both categories contain identical medical images. Closer inspection of the Fig. 6d and Tab. II reveals that SvANet is the only method that achieved a “V” trend in the ATLAS dataset, with the best mDice of 89.79% for segmenting ultra-small compared to small and all sizes of livers and tumors, underscoring SvANet’s capability to effectively discriminate ultra-small medical objects. Additionally, SvANet consistently displays narrow error bars (shown as color bands) across three object scales, indicating its robustness in accurately recognizing medical objects.
IV-C Results for the Dataset with Only Ultra-Small Objects
To further evaluate the performance of SvANet in distinguishing ultra-small medical objects, experiments were conducted in the SpermHealth dataset, which exclusively has sperms with an area ratio of less than 1%. As shown in Tab. III, SvANet secured top performance in sperm segmentation within the SpermHealth dataset, achieving 72.58% in mDice, 61.44% in mIoU, 13.06×10-4 in MAE, 72.50% in sensitivity, and 72.51% F2 score. SvANet’s performance in sperm segmentation notably exceeded that of other models, surpassing them by up to 15.88% in F2 score, 14.99% in sensitivity, 14.64% in mDice, and 11.86% in mIoU. Besides, the performance metrics (mDice, mIoU, sensitivity, and F2 score) gained in the SpermHealth dataset are significantly lower than those observed in ISIC 2018, PolypGen, ATLAS, and KiTS23 for all tested models, with a gap of $>$10%, because all sperms have an area lower than 1%, presenting limited learnable features and posing more significant challenges for differentiation.
@@ IV-C Results for the Dataset with Only Ultra-Small Objects
To quantify the robustness and adaptability of SvANet vs. other SOTA methods, receiver operating characteristic (ROC) curves of tested methods in the SpermHealth dataset are employed and illustrated in Fig. 7. The ROC curve of SvANet, represented by the red line in Fig. 7, blends nearest towards the top-left corner, with the highest area under the curve (AUC) of 0.9905, surpassing other SOTA methods by up to AUC of +0.008. Besides, the ROC curves of UNet++ is close to the lower-right corner and under all other curves, with the lowest AUC of 0.9825. The ROC and AUC results of UNet++ are consistent with Tab. III, demonstrating that UNet++ struggled to recognize sperms.
IV-D Ablation Studies
Unless otherwise specified, all ablation studies were conducted in the SpermHealth dataset for the sake of simplicity.
IV-D1 Main Components Ablation
To investigate the influence of each core module of SvANet (i.e., MCBottleneck, MCAttn, cross-scale guidance, SvAttn, and AssemFormer), ablation studies were conducted and discussed in this section. As indicated in the first and second columns of Tab. IV, the inclusion of MCBottleneck and MCAttn results in improvements of +0.15% & +0.52% mDice and +0.40% & +0.22% sensitivity, respectively, suggesting that both modules significantly contribute to enhancing the diagnosis and accuracy of positive cases. Furthermore, when equipped with cross-scale guidance and SvAttn, SvANet achieves an additional +0.41% & +0.24% in mDice and +0.53% & +0.51% in sensitivity. The ablation of AssemFormer leads to an increase of +0.16% in mDice and +0.67% in sensitivity. The p-value for mDice is below 0.05, confirming the result’s reliability. In summary, MCAttn and cross-scale guidance are critical modules for enhancing prediction accuracy. At the same time, AssemFormer and MCBottleneck are vital in improving the accuracy of positive diagnostics in medical object segmentation.
TABLE IV: Ablation studies results on the main components of SvANet. The best results are underlined in bold.